# Mind the gap: Distinguishing disc substructures and their impact on the composition of the inner disc

### v1

@@ Acknowledgements.
We thank the anonymous reviewer for the insightful comments which helped to improve the clarity of this paper. The authors acknowledge the support of the DFG priority program SPP 1992 “Exploring the Diversity of Extrasolar Planets” (BI 1880/3-1).References
@@
- Pinilla et al. (2024) Pinilla, P., Benisty, M., Waters, R., Bae, J., & Facchini, S. 2024, arXiv e-prints, arXiv:2403.07057
@@ A.1 Gas evolution
where is a dimensionless parameter that describes the strength of turbulence (Shakura & Sunyaev 1973), is the Keplerian angular velocity, and is the isothermal sound speed which is in turn related to the disc’s midplane temperature following
@@ Appendix D Evolution of inner disc composition as a function of gap formation time
Here we show the plots illustrating how the gap formation time influences the time evolution of the C/O ratio (Fig. 6), the O/H ratio (Fig. 7) and the C/H ratio (Fig. 8) in the inner disc.
@@ Appendix E Connecting gap depth to planet mass
where is related to the properties of the planet and the disc following (Crida et al. 2006)
@@
In Fig. 9 we show how Eq. 16 changes with at and 10 AU for different disc viscosities. The gap depth increases ( decreases) with increasing planet mass, as expected. More importantly, the plot also shows that we need more massive planets to achieve the same value of gap depth when (1) the disc viscosity increases, or (2) the planet is located further out in the disc. This is related to the gas scale height which increases with disc viscosity and orbital distance (Bitsch et al. 2015; Savvidou et al. 2020).

### v2

@@ Acknowledgements.
We thank the anonymous reviewer for the insightful comments which helped to improve the clarity of this Letter. The authors acknowledge the support of the DFG priority program SPP 1992 “Exploring the Diversity of Extrasolar Planets” (BI 1880/3-1).References
@@
- Pinilla et al. (2024) Pinilla, P., Benisty, M., Waters, R., Bae, J., & Facchini, S. 2024, A&A, 686, A135
@@ A.1 Gas evolution
where is a dimensionless parameter that describes the strength of turbulence (Shakura & Sunyaev 1973), is the Keplerian angular velocity, and is the isothermal sound speed which is in turn related to the disc’s midplane temperature following:
@@ Appendix D Evolution of inner disc composition as a function of gap formation time
Here, we show the plots illustrating how the gap formation time influences the time evolution of the water abundance (Fig. 5), the C/O ratio (Fig. 6), the O/H ratio (Fig. 7), and the C/H ratio (Fig. 8) in the inner disc.
@@ Appendix E Connecting gap depth to planet mass
where is related to the properties of the planet and the disc following (Crida et al. 2006):
@@
In Fig. 9, we show how Eq. 16 changes with at and 10 AU for different disc viscosities. The gap depth increases ( decreases) with increasing planet mass, as expected. More importantly, the plot also shows that we need more massive planets to achieve the same value of gap depth when: (1) the disc viscosity increases or (2) the planet is located further out in the disc. This is related to the gas scale height which increases with disc viscosity and orbital distance (Bitsch et al. 2015; Savvidou et al. 2020).